# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ B.2 Annealed Langevin Dynamics in Sub-Gaussian Mixtures
where ${\mathbf{R}}\in{\mathbb{R}}^{d\times r}$ an orthonormal basis of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$ and ${\mathbf{N}}\in{\mathbb{R}}^{d\times n}$ an orthonormal basis of the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$. Now, we prove Theorem 4 by applying the techniques developed in Appendix A.1 and A.3 via substituting ${\nu}^{2}$ and $\frac{{\nu}^{2}}{1-c_{{\nu}}}$ with $\frac{{\nu}^{2}+{\sigma}_{t}^{2}}{1-c_{{\nu}}}$ at time step $t$. Note that for all $t\in\left\{{0}\right\}\cup[T]$, Assumption 3.iv. implies ${\nu}_{0}^{2}+{\sigma}_{t}^{2}>\max\left\{{1,\frac{4(c_{L}^{2}+c_{{\nu}}c_{L})}{c_{{\nu}}(1-c_{{\nu}})}}\right\}\frac{{\nu}_{\max}^{2}+{\sigma}_{t}^{2}}{1-c_{{\nu}}}$ because $c_{{\sigma}}\geq{\sigma}_{t}$.
First, by Proposition 2, suppose that the sample is initialized in the distribution $P^{(0)}_{{\sigma}_{0}}$, then with probability at least $1-\exp(-\Omega(d))$, we have
$\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3({\nu}_{0}^{2}+{\sigma}_{0}^{2})}{4}+\frac{{\nu}_{\max}^{2}+{\sigma}_{0}^{2}}{4(1-c_{{\nu}})}\right)d.$ (21)
Then, with the assumption that the initialization satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3({\nu}_{0}^{2}+{\sigma}_{0}^{2})}{4}+\frac{{\nu}_{\max}^{2}+{\sigma}_{0}^{2}}{4(1-c_{{\nu}})}\right)d$, the following proposition similar to Proposition 4 shows that $\left\|{\mathbf{n}_{t}}\right\|$ remains large with high probability.
Proposition 7.
Consider a distribution $P$ satisfying Assumption 3. We follow annealed Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps with noise level $c_{{\sigma}}\geq{\sigma}_{0}\geq{\sigma}_{1}\geq\cdots\geq{\sigma}_{T}\geq 0$ for some constant $c_{{\sigma}}>0$. Suppose that the initial sample satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3({\nu}_{0}^{2}+{\sigma}_{0}^{2})}{4}+\frac{{\nu}_{\max}^{2}+{\sigma}_{0}^{2}}{4(1-c_{{\nu}})}\right)d$, then with probability at least $1-T\cdot\exp(-\Omega(d))$, we have that $\left\|{\mathbf{n}_{t}}\right\|^{2}>\left(\frac{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}{2}+\frac{{\nu}_{\max}^{2}+{\sigma}_{t}^{2}}{2(1-c_{{\nu}})}\right)d$ for all $t\in\{0\}\cup[T]$.
Proof of Proposition 7.
We prove Proposition 7 by induction. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:

### v3

@@ B.2 Proof of Theorem 6
where ${\mathbf{R}}\in{\mathbb{R}}^{d\times r}$ an orthonormal basis of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$ and ${\mathbf{N}}\in{\mathbb{R}}^{d\times n}$ an orthonormal basis of the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$. Now, we prove Theorem 6 by applying the techniques developed in Appendix A.2 and B.1 via substituting ${\nu}^{2}$ and $\frac{{\nu}^{2}}{1-c_{{\nu}}}$ with $\frac{{\nu}^{2}+{\sigma}_{t}^{2}}{1-c_{{\nu}}}$ at time step $t$. Note that for all $t\in\left\{{0}\right\}\cup[T]$, Assumption 2.iv. implies ${\nu}_{0}^{2}+{\sigma}_{t}^{2}>\max\left\{{1,\frac{4(c_{L}^{2}+c_{{\nu}}c_{L})}{c_{{\nu}}(1-c_{{\nu}})}}\right\}\frac{{\nu}_{\max}^{2}+{\sigma}_{t}^{2}}{1-c_{{\nu}}}$ because $c_{{\sigma}}\geq{\sigma}_{t}$.
Then, with the assumption that the initialization satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\left(\frac{3({\nu}_{0}^{2}+{\sigma}_{0}^{2})}{4}+\frac{{\nu}_{\max}^{2}+{\sigma}_{0}^{2}}{4(1-c_{{\nu}})}\right)d$, we prove Theorem 6 via showing that
${\mathbb{P}}\left(\forall t\in[T],\,\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\left(\frac{{\nu}_{0}^{2}+{\sigma}_{t}^{2}}{2}+\frac{{\nu}_{\max}^{2}+{\sigma}_{t}^{2}}{2(1-c_{{\nu}})}\right)d\right)\geq 1-T\cdot\exp\left(-\Omega(d)\right).$
Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases: